# 11email: {giovanni.catalani}[at]airbus.com

### v1

@@ Abstract
Recent advances in Neural Fields have enabled powerful, discretization-invariant methods for learning neural operators that approximate solutions of Partial Differential Equations (PDEs) on general geometries. Building on these developments, we introduce enf2enf, an encoder–decoder methodology for predicting steady-state Partial Differential Equations with non-parameterized geometric variability, based on recently proposed Equivariant Neural Field architectures. In enf2enf, input geometries are encoded into latent point cloud embeddings that inherently preserve geometric grounding and capture local phenomena. The resulting representations are then combined with global parameters and directly decoded into continuous output fields, thus efficiently modeling the coupling between geometry and physics. By leveraging the inductive biases of locality and translation invariance, our approach is able to capture fine-scale physical features as well as complex shape variations, thereby enhancing generalization and physical compliance. Extensive experiments on a high-fidelity aerodynamic dataset, a hyper-elastic material benchmark, and multi-element airfoil geometries, demonstrate that the proposed model achieves superior or competitive performance compared to state-of-the-art graph based, operator learning, and neural field methods. Notably, our method supports real time inference and zero-shot super-resolution, enabling efficient training on low-resolution meshes while maintaining high accuracy on full-scale discretizations.
Keywords:
Partial Differential Equations, Neural Operators, Neural Fields
@@ 1 Introduction
Partial Differential Equations (PDEs) are fundamental to modeling physical phenomena across numerous scientific and engineering domains [11, 37]. They are essential in fluid dynamics, heat transfer, structural mechanics, electromagnetics, and quantum mechanics mathematical modeling. In aerospace engineering particularly, Computational Fluid Dynamics (CFD) has become an indispensable tool for aircraft design, enabling high-fidelity simulations of complex aerodynamic phenomena. These simulations, primarily based on numerical approximations of the Reynolds Averaged Navier Stokes (RANS) equations [29], provide detailed predictions of flow fields, aerodynamic performance, or structural loads. However, the computational cost of traditional numerical methods remains a significant bottleneck [13]. The complexity arises from the large number of degrees of freedom required to capture relevant physics, mesh refinement requirements for accurate boundary layer resolution, and iterative solution procedures. Additionally, in typical engineering applications, multiple solutions of the same governing equations are required to extensively explore the design space due to variations in domain geometry (boundary conditions) or input conditions, such as fluid density or inflow velocity in the case of aerodynamics. The need for less computationally demanding simulations and the abundance of available data motivate the development of data-driven PDE surrogate models, that can offer remarkable speedups compared to traditional numerical methods, while reducing the level of fidelity [7].
Traditional surrogate models based on modal decomposition such as Proper Orthogonal Decomposition [4] (POD or Principal Component Analysis) coupled with interpolation techniques [31] (Gaussian Process Regression, Radial Basis Functions) excel on fixed geometries and in modeling linear phenomena, but it is not trivial to extend them to handle changing geometries [8] and levels of discretization. Deep neural networks, grounded in universal approximation theory, have demonstrated the ability to represent highly complex functions in high-dimensional spaces, by extracting meaningful representations from complex data [3, 6].
This representational power justifies successes in areas such as image processing [22] and generative modeling, and can extend to the area of PDE modeling, which concerns learning high-dimensional input–output mappings and capturing multiscale phenomena across complex solution manifolds.
Crucially, Deep Learning architectures leverage certain assumptions about the solution space, referred to as inductive biases [16], and restrict the function class to those that align with the underlying dataset or physical constraints, thereby promoting better generalization beyond the training set. Convolutional Neural Networks (CNNs) [21], employ convolution operations enforcing spatial locality and translation invariance. This inductive bias has been shown to be particularly effective for PDE-related tasks [1], where local interactions (e.g., localized regions of fluid flow or stress concentration) are detected in different absolute locations in the spatial domain. The reliance of CNN on uniform grids (or coupling with an interpolation routine) renders them unsuitable for physical applications that employ unstructured meshes [10]. Graph Neural Networks [6] (GNNs) address this issue by extending local message passing to irregular graph topologies, thus accommodating arbitrary mesh structures [28]. However, GNN based models typically encode the mesh topology and resolution directly into their representation, leading to biases when generalizing to unseen discretizations [14] or mesh topologies. Moreover, the requirement to store node neighborhoods and use an increasing number of message passing layers can result in substantial memory and computational overhead, especially for large scale industrial meshes with millions of nodes [18]. These factors hinder generalization across varying mesh resolutions and geometries, emphasizing the need for discretization invariant approaches that can efficiently handle large scale PDEs [25, 9]. Operator Learning aims to directly approximate the PDE solution operator, mapping the input (e.g., boundary conditions, source terms, or geometrical parameters) and output function spaces [2]. By treating both inputs and outputs as functions, it is possible to obtain continuous representations of physical fields that can be queried at arbitrary points in the spatial domain. DeepONet [26] and the Fourier Neural Operator (FNO) [24], are two prominent approaches that have demonstrated compelling results in various PDE modeling tasks. It is not trivial to extend these architectures to handle non-uniformly sampled grids [23] (FFT can only be computed efficiently on structured grids) and unparameterized geometric variations, such as in the case of geometry aware inference, for instance for the predictions of the pressure fields over airfoils.
Conditional Neural Fields have emerged as a powerful framework to construct Neural Operator architectures [38] offering several attractive properties including discretization independence and good spectral convergence thanks to Fourier Feature Networks [36] or sinusoidal [34] layer activations. CORAL [32] encodes latent representations of physical fields (referred to as functa) using a meta-learning approach [12]. For geometry aware inference, two separate encoders are used to encode functa for input functions (geometry) and output physical fields, which are then separately mapped in the latent space: this approach, referred to as encode-process-decode, has also been extended for 3D applications concerning aircraft aerodynamic simulations [9]. Neural Field approaches that rely on a global shared network and latent vector modulation often lack a strong inductive bias in both the neural field architecture and the latent space structure. Because each physical field is encoded as a single global vector, the representation lacks grounding in geometry [39]. This results in latent representations with limited downstream utility, making it challenging to model the effect of local geometric features and large scale physical phenomena[27]. Moreover, existing methods (based on encode-process-decode) typically minimize an error metric on the learned latent space (when mapping input and output latent spaces) rather than directly penalizing discrepancies in the physical output fields themselves, leading to potential mismatches between the latent space objective and the final PDE solution accuracy. Such limitations often necessitate strong regularization on the latent representations to mitigate overfitting [12, 27].
Recent work on Equivariant Neural Fields (ENFs) [39] has shown how incorporating geometric structure in the latent space can significantly improve representation learning and downstream tasks. By encoding fields into latent point clouds that are grounded in the spatial domain, local conditioning and relevant symmetries can be enforced, leading to improved sample efficiency. Locality priors, in particular, are important for PDE modeling (i.e. the state at a given point is predominantly influenced by nearby geometric and flow conditions) and several works have exploited attention mechanisms [33, 40] to implement local conditioning, demonstrating enhanced generalization and reduced computational overhead by focusing the network’s capacity on the most relevant regions of interest.
Accurately modeling geometric variability is critical for predicting physical phenomena: variations of a wing shape or of a structural component directly influence the resulting flow-field or stress distribution, with important consequences for industrial design optimization. Equivariant Neural Fields have been demonstrated for time-dependent PDE modeling on fixed geometries where each time step is encoded into a geometrically grounded latent representation and evolved by a separate latent time-processing module [20]. In contrast, our work addresses a different challenge: modeling variable geometries (and operating conditions) to directly infer the resulting steady-state physical fields. To achieve this, we propose enf2enf, a novel geometry aware operator learning framework that leverages ENFs representations to tackle steady state PDEs on complex geometries. In our encoder-decoder framework, the input geometric fields (such as Signed Distance Functions) are encoded into latent representations that are grounded in the physical domain and are directly mapped to the output fields by a downstream decoder, which handles the processing of both local (geometric) and global (inflow) conditioning. Crucially, compared to previous approaches based on encode-process-decode pipeline [32, 9], we bypass the separate latent space processor which typically introduces an auxiliary latent objective decoupled from the final PDE reconstruction error. This leads to a simpler architecture, reducing computational overhead in training and inference, and yields a more direct interaction between geometry and physics, potentially enhancing generalization performance. Furthermore, we demonstrate competitive performance on a material modeling benchmark and a large scale aerodynamic dataset, characterized by high-resolution CFD simulations over complex airfoil geometries. Our extensive baseline comparisons include state-of-the-art operator learning frameworks, GNN-based methods, and conventional neural field approaches. Detailed analyses reveal that the added inductive biases such as localized geometric encoding and translation equivariant attributes enhance the model’s ability to generalize across varying mesh resolutions and to capture both global inflow conditions and localized phenomena with improved accuracy and physical consistency.
@@ 2 Methodology
Encoder Training
Decoder Training
@@
Figure 1: Architecture overview of enf2enf showing different computational paths during training and inference. The input field is processed by encoder $\mathcal{E}_{\text{a}}$ to produce latent variables $z_{a}$ defined as point clouds (green) with node features $c_{i}^{\text{in}}$, which are nodewise concatenated ($\oplus$) with parameters $\mu$ (leading to output latents $z_{u}$ in red) before being decoded by $\mathcal{D}_{\text{u}}$ to generate the output field.
@@ 2.1 Problem Formulation
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain, where $d$ is the spatial dimension. We define two separable Banach spaces $\mathcal{A}=\mathcal{A}(\Omega;\mathbb{R}^{d_{a}})$ and $\mathcal{U}=\mathcal{U}(\Omega;\mathbb{R}^{d_{u}})$, consisting of functions mapping from $\Omega$ to $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$ respectively. The input function space $\mathcal{A}(\Omega)$ is defined as:
$\mathcal{A}(\Omega)=\mathcal{A}_{\text{geom}}(\Omega;\mathbb{R})\times\mathcal{A}_{\mu}(\Omega;\mathbb{R}^{d_{\mu}}),$ (1)
where $a_{\text{geom}}\in\mathcal{A}_{\text{geom}}$ encodes the geometry (e.g., Signed Distance Functions continuously describing the shape boundaries), and $\mu\in\mathcal{A}_{\mu}$ represents global parameters (e.g., inflow conditions, material properties). The space $\mathcal{U}$ contains functions mapping to physical field solutions, such as velocity, pressure or stress distribution, that satisfy the governing partial differential equations under the conditions specified by elements of $\mathcal{A}$. The governing partial differential equations define a nonlinear operator mapping between these spaces:
$\mathcal{G}^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}.$ (2)
We seek to approximate $\mathcal{G}^{\dagger}$ with a parameterized operator $\mathcal{G}_{\theta}:\mathcal{A}\rightarrow\mathcal{U}$, where $\theta\in\Theta$ represents the learnable parameterization. Given a probability measure $\nu$ on $\mathcal{A}$, we aim to minimize the approximation error in the continuous setting:
[…]
We seek to learn this operator from a collection of input-output pairs $(a_{i},u_{i})\in\mathcal{A}\times\mathcal{U}$ (where again $a_{i}=(a_{geom,i},\mu_{i})$) such that $u_{i}=\mathcal{G}^{\dagger}(a_{i})$. Due to the numerical nature of the solutions, at training time we have access to $n_{tr}$ pairs of input and output functions $(a_{i},u_{i})_{i=1}^{n_{tr}}$ evaluated on unstructured grids. Specifically, each sample is observed on its own computational mesh $\mathcal{X}_{i}$. We denote by $a|_{\mathcal{X}_{i}}=(a(x))_{x\in\mathcal{X}_{i}}$ and $u|_{\mathcal{X}_{i}}=(u(x))_{x\in\mathcal{X}_{i}}$ the vectors of function values evaluated over the sample mesh. This leads to the empirical minimization problem:
$\theta^{\dagger}=\arg\min_{\theta\in\Theta}\frac{1}{n_{tr}}\sum_{i=1}^{n_{tr}}\sum_{x\in\mathcal{X}_{i}}\|u_{i}(x)-\mathcal{G}_{\theta}(a_{i})(x)\|^{2}.$ (4)
@@ 2.2 Model Architecture and Training
Our architecture uses two main components: $f_{\theta_{a},z_{a}}$, $f_{\theta_{u},z_{u}}$ for input and output functions respectively, both implemented as Equivariant Neural Fields (ENFs) [39], where $\theta_{a},\theta_{u}$ are global shared parameters and $z_{a},z_{u}$ are sample-specific latent point cloud representations of input and output functions. Input functions $a_{i}$ are encoded into latent representations using the input encoder $z_{a_{i}}=\mathcal{E}_{\text{a}}({a_{i}})$ by solving an optimization problem $f_{\theta_{a},z_{a_{i}}}(x)|_{\mathcal{X}_{i}}=a_{i}$. This procedure, known as auto-decoding, is better detailed in the following dedicated paragraph. Obtained input representations are then (nodewise) concatenated with global parameters $\mu$ (when present), and decoded to output physical fields through the output spatial decoder $\mathcal{D}_{u}$:
$\hat{u}_{i}(x)=D_{u}(z_{u_{i}})(x)=f_{\theta_{u},z_{u_{i}}}(x),\quad z_{u_{i}}=\text{cat}\big{(}z_{a_{i}},\,\mu\big{)}.$ (5)
The resulting operator $\mathcal{G}_{\theta}$ can then be expressed as the composition:
$\mathcal{G}_{\theta}=\mathcal{D}_{\text{u}}\circ\mathcal{E}_{\text{a}}.$ (6)
@@
Training proceeds in two stages, sketched in Figure 1, by minimizing the reconstruction losses for both input and output fields:
$\mathcal{L}^{a}=\sum_{i=1}^{n_{tr}}\|a_{i}-\mathcal{D}_{\text{a}}(\mathcal{E}_{\text{a}}(a_{i}))\|^{2}_{\mathcal{A}},\quad\mathcal{L}^{u}=\sum_{i=1}^{n_{tr}}\|u_{i}-\mathcal{D}_{\text{u}}(\mathcal{E}_{\text{a}}(a_{i}))\|^{2}_{\mathcal{U}}$ (7)
where the input decoder component $D_{a}$ is used for learning to reconstruct the input functions from input latents, and it is not involved in the inference process. In the following we describe the specific architecture of input and output ENFs, and the encoding process.
2.2.1 Translation Equivariant Neural Fields.
Equivariant Neural Fields (ENFs) are coordinate-based neural networks that can be equipped with different symmetries through the choice of bi-invariant attributes [39]. These attributes define the relationship between spatial coordinates and latent representations in a way that preserves the desired symmetries and invariances of the problem. Compared to standard conditional Neural Fields autodecoder architectures that extract sample-wise representations in the form of 1D arrays, ENFs encode functions on latent point cloud representations embedded in the physical domain and as such grounded in geometry [20]:
$z:=\{(p_{i},c_{i})\}_{i=1,..,N_{lat}}$ (8)
where the latent features $c_{i}\in\mathbb{R}^{l_{d}}$ are localized in the latent positions $p_{i}\in\mathbb{R}^{d}$ of the physical domain. For fluid flow problems and general physical PDEs without evident rotational or scaling symmetries, translation equivariance represents a strong inductive bias to characterize the physical system. Many fundamental physical phenomena, such as vortices or shocks, can occur at different absolute locations while maintaining identical local behavior and properties. Therefore, we employ relative positions as our bi-invariant attribute:
$b_{i,m}=\gamma(x_{m}-p_{i})$ (9)
where $p_{i}$ represents the position of the $i$-th latent point and $x_{m}$ is the query position. Fourier Feature Encoding $\gamma$ is used on the inputs, as in standard Neural Fields to overcome the spectral bias and capture the high frequency components of the target signals. The output of the neural field architecture at a specific location $x_{m}$ is obtained as the weighted sum of the different contributions from the latent points:
$f_{\theta_{a}}(x_{m};z)=W_{o}\sum_{i=1}^{N_{lat}}\text{att}_{m,i}v(b_{i,m},c_{i})$ (10)
where the attention weights are given by:
$\text{att}_{m,:}=\text{softmax}_{i=1,...,N_{lat}}\left(\frac{q(b_{i,m})^{T}k(c_{i})}{\sqrt{d_{k}}}-\sigma||x_{m}-p_{i}||^{2}\right)$ (11)
and the second term (whose magnitude is controlled by the gaussian window size scalar $\sigma$) penalizes the distance of the query point from the latent position, therefore enhancing local conditioning. The query and key functions are parameterized as:
$q(b_{i,m})$ $=W_{q}b_{i,m}$ (12)
$k(c_{i})$ $=W_{k}c_{i}$ (13)
while the value is obtained by linear transformation of the bi-invariants through scale and shift modulation:
$v(b_{i,m},c_{i})=(W_{v}c_{i})\odot(W_{a}^{\gamma}b_{i,m})+W_{a}^{\beta}b_{i,m}.$ (14)
2.2.2 Encoding
For each input function $a$ (geometric field), the encoding process follows a meta-learning optimization procedure based on CAVIA [41]. Specifically, given the shared network parameters $\theta_{a}$, the latent point-cloud representations $z_{a}=\{(p_{i},c_{i})\}_{i=1,...,N_{lat}}$ are obtained by solving an optimization problem (omitting the index $i$ for brevity):
$c^{(0)}$ $=0\quad p^{(0)}=p^{(K)}$ (15)
$c^{(k+1)}$ $=c^{(k)}-\alpha\nabla_{c^{(k)}}\mathcal{L}^{a}_{\mathcal{X}_{i}}(f_{\theta_{a},c^{(k)}},a)\quad\text{for }0\leq k\leq K-1.$ (16)
where compared to [39, 20], we initialize the latent positions $p$ for all samples to cover uniformly the input domain and decide not to optimize them, as the updating their value results in less stable training results. Here, $\alpha$ is the inner loop learning rate, $K$ is the number of optimization steps (typically set to 3), and $\mathcal{L}^{a}_{\mathcal{X}_{i}}$ is the reconstruction loss evaluated over the training grid defined on the input domain. During training of the input encoder, the optimization problem (inner loop) is solved in parallel with the global parameters $\theta_{a}$ optimization (outer loop). At inference, the global parameters are fixed, as it is sufficient to run the inner loop in few gradient descent steps ($K$) to encode a new geometry.

### v2

# NeurIPS 2025 AI for Science workshop

@@ Abstract
Advances in neural operators have introduced discretization invariant surrogate models for PDEs on general geometries, yet many approaches struggle to encode local geometric structure and variable domains efficiently. We introduce enf2enf, a neural field approach for predicting steady-state PDEs with geometric variability. Our method encodes geometries into latent features anchored at specific spatial locations, preserving locality throughout the network. These local representations are combined with global parameters and decoded to continuous physical fields, enabling effective modeling of complex shape variations. Experiments on aerodynamic and structural benchmarks demonstrate competitive or superior performance compared to graph-based, neural operator, and recent neural field methods, with real-time inference and efficient scaling to high-resolution meshes.
@@ 1 Introduction
Partial Differential Equations (PDEs) are fundamental to modeling physical phenomena across numerous scientific and engineering domains [1, 2]. Computational Fluid Dynamics (CFD), for instance, has become an indispensable tool for aircraft and vehicle design, enabling high-fidelity simulations of complex aerodynamic phenomena. However, the computational cost of traditional numerical methods remains a significant bottleneck [3]. The complexity arises from the large number of degrees of freedom required to capture multiscale physics, mesh refinement requirements for accurate boundary layer resolution, and iterative solution procedures. The need for computationally efficient simulations and the abundance of available data motivate the development of data-driven PDE surrogate models that can offer remarkable speedups compared to traditional numerical methods, while reducing the level of fidelity [4].
In this context, deep neural networks have complemented existing reduced order models based on linear dimensionality reduction and interpolation [5], which are typically limited to handle fixed geometries, providing significant advantages for physical modeling tasks with geometric variability. Early neural network architectures for PDE modeling exploited spatial inductive biases through Convolutional Neural Networks (CNNs) [6] on regular grids or Graph Neural Networks (GNNs) [7] on irregular meshes. However, CNNs necessitate costly interpolation procedures when handling unstructured computational grids [8], while many GNN-based approaches encounter significant computational bottlenecks on large-scale industrial meshes [9] and demonstrate poor generalization across different mesh resolutions [10]. To overcome these discretization-dependent limitations, Neural operators have emerged as a promising alternative by framing the problem as learning mappings between infinite-dimensional function spaces rather than learning input-output relationships defined on fixed resolutions or specific discretizations. This paradigm enables discretization-invariant learning across different mesh resolutions. Notable examples include DeepONet [11] and Fourier Neural Operators (FNO) [12], which have shown impressive performance across various PDE benchmarks. Nevertheless, these methods suffer from architectural constraints: DeepONet requires manual design of input function encodings, while FNO is fundamentally limited to structured grids [13] and periodic boundary conditions, hindering their deployment in complex geometric scenarios. More recent approaches based on Neural Fields, such as CORAL [14] address these limitations by encoding input functions into compact global latent representations using implicit neural networks. However, these approaches rely on global latent representations that struggle to preserve local geometric structure. This limitation is particularly problematic for PDEs on variable geometries, where spatially localized physical phenomena are crucial for accurate modeling. The idea of using spatially grounded latent representation has been recently introduced for dynamics modeling on fixed domains [15] using Equivariant Neural Field. However, the challenge of geometry-aware inference for steady-state PDEs across changing geometries has not been tackled, despite its paramount importance in practical engineering scenarios.
Contribution.
To address these limitations, we propose enf2enf, a novel geometry-aware operator learning framework specifically designed for steady-state PDEs on variable geometries. Our approach employs a geometry encoder that extracts spatially-grounded point cloud embeddings from input geometric fields (such as Signed Distance Functions), such that each latent point is anchored to a specific spatial location and encodes local geometric features within its neighborhood. The decoder then processes these local geometric encodings by combining them with global operating conditions (such as inflow parameters). Self-attention mechanisms facilitate message passing between latent points, enabling spatial communication. This encoder-decoder design bypasses the encode-process-decode pipeline employed in previous approaches [14], eliminating the intermediate latent space processor that introduces auxiliary objectives decoupled from the final PDE reconstruction error. The experimental validation focuses on applied PDE benchmarks, including structural modeling and large-scale aerodynamic simulations.
2 Related Work
Mesh-based surrogate models.
Traditional surrogate modeling approaches for PDEs rely on modal decomposition techniques such as Proper Orthogonal Decomposition (POD) [5] combined with Gaussian Process Regression (GPR) for interpolation [16]. While effective on fixed geometries, extending these methods to handle geometric variability remains challenging. Mesh morphing variants (MMGP) [17] attempt to address this through reference mesh deformation but are constrained to similar topologies. Graph Neural Networks (GNNs), particularly MeshGraphNets [7], emerged as alternatives for irregular meshes but face computational bottlenecks on large-scale problems and exhibit poor generalization across different mesh resolutions [10].
Neural operators.
Operator learning addresses these limitations by modeling mappings between infinite-dimensional function spaces rather than discrete mesh representations. DeepONet [11] enables querying at arbitrary coordinates but requires input functions to be observed on predefined grids, necessitating identical observation patterns across training and testing. Fourier Neural Operators (FNO) [12] leverage fast Fourier transforms for efficient spectral domain computations but are fundamentally restricted to structured grids. Extensions like Geo-FNO [13] and Geometry-Informed Neural Operators (GINO) attempt to handle irregular geometries but do not show remarkable gains in performance, while lacking explicit mechanisms to encode geometric variability. Recent transformer-based approaches like Transolver [18] incorporate attention mechanisms for spatial reasoning, and are able to locally extract slices, thereby improving accuracy for the surrogate modeling task on variable geometries.
Neural fields or Implicit Neural Representations.
Implicit Neural Representations (INRs) model spatial data as continuous functions [19, 20], enabling coordinate-based querying at arbitrary locations. CORAL [14] employs meta-learning to encode input functions into global latent vectors, then processes these representations in latent space before decoding to output fields through an encode-process-decode pipeline. However, compressing entire geometric configurations into single global vectors discards spatial locality and geometric structure. Equivariant Neural Fields (ENFs) [21] address this limitation by introducing geometrically grounded latent representations, but existing applications focus exclusively on time-dependent PDE modeling with fixed geometries [15], still employing an intermediate processor network and without incorporation of global operating conditions. In contrast, our approach employs ENFs solely as geometry encoders, eliminates the intermediate processor, and directly combines local geometric conditioning with global operating parameters within the decoder network.
@@ 3 Methodology
Input Enc. Training
Output Dec. Training
@@
Figure 1: Overview of the enf2enf architecture. The input geometry field $a_{i}$ is encoded by $\mathcal{E}_{\mathrm{a}}$ into a spatially anchored latent point cloud $z_{a}=\{(p_{j},c_{j}^{\text{in}})\}_{j=1}^{N_{\text{lat}}}$ (green). Each latent feature $c_{j}$ is node-wise concatenated with the global parameters $\mu$ to yield the conditioned set $z_{u}$ (red), which the decoder $\mathcal{D}_{\mathrm{u}}$ maps to the predicted physical field $\hat{u}_{i}$ via a composition of latent self attention (SA) and equivariant cross attention (ECA) decoding. The different training and inference paths are highlighted with the respective losses.
@@ 3.1 Problem Formulation
We consider the operator learning problem for steady state PDEs on variable geometries. Given input functions $a_{i}$ consisting of geometric descriptors (e.g., signed distance functions) and global parameters $\mu_{i}$ (e.g., inflow conditions, material properties), we seek to learn the mapping to output physical fields $u_{i}$ (e.g., pressure, stress) that satisfy the governing PDEs.
Formally, let $\Omega\subset\mathbb{R}^{d}$ be a bounded spatial domain where $d\in\{2,3\}$. We define the input space as $\mathcal{A}=\mathcal{A}_{\text{geom}}(\Omega)\times\mathcal{A}_{\mu}$, where $a_{\text{geom}}:\Omega\to\mathbb{R}^{n_{a}}$ encodes the geometry as functions over $\Omega$ (with $n_{a}$ denoting the input field dimension, e.g., $n_{a}=1$ for scalar SDF) and $\mu\in\mathcal{A}_{\mu}\subset\mathbb{R}^{l_{\mu}}$ represents global parameters. The governing PDEs define a nonlinear operator $\mathcal{G}^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}(\Omega)$ mapping to the solution function space $\mathcal{U}(\Omega)$, where $u:\Omega\to\mathbb{R}^{n_{u}}$ (with $n_{u}$ denoting the output field dimension).
We approximate $\mathcal{G}^{\dagger}$ with a parameterized operator $\mathcal{G}_{\theta}$ learned from training pairs $(a_{i},u_{i})$ observed on unstructured meshes $\mathcal{X}_{i}\subset\Omega$:
$\theta^{\dagger}=\arg\min_{\theta}\frac{1}{n_{tr}}\sum_{i=1}^{n_{tr}}\sum_{x\in\mathcal{X}_{i}}\|u_{i}(x)-\mathcal{G}_{\theta}(a_{i})(x)\|^{2}.$ (1)
We implement the neural operator architecture $\mathcal{G}_{\theta}=\mathcal{D}_{\text{u}}\circ\mathcal{E}_{\text{a}}$ where:
  • The encoder $\mathcal{E}_{\text{a}}$ maps geometric inputs $a_{\text{geom}}$ to spatially-grounded representations $z_{a}$.
  • The decoder $\mathcal{D}_{\text{u}}$ combines these local geometric features with global parameters $\mu$ to produce output fields $u$.
@@ 3.2 Model Architecture and Training
Our architecture uses two main components: $f_{\theta_{a},z_{a}}:\mathbb{R}^{d}\to\mathbb{R}^{n_{a}}$, $f_{\theta_{u},z_{u}}:\mathbb{R}^{d}\to\mathbb{R}^{n_{u}}$ for input and output functions respectively. Both are implemented as conditional neural field architectures, where $\theta_{a},\theta_{u}$ are global shared parameters and $z_{a},z_{u}$ are sample-specific latent point cloud representations:
$z=\{(p_{j},c_{j})\}_{j=1,..,N_{lat}}$
where latent features $c_{j}\in\mathbb{R}^{l_{d}}$ are localized at spatial positions $p_{j}\in\mathbb{R}^{d}$ within the physical domain.
Input functions $a_{i}$, typically representing the geometry (such as SDF fields), are encoded into latent representations using the input encoder $z_{a_{i}}=\mathcal{E}_{\text{a}}({a_{i}})$ via meta-learning [22], detailed below. Obtained local input representations are then processed with global parameters $\mu$ (when present), and decoded to output physical fields through the output spatial decoder $\mathcal{D}_{u}$:
$\hat{u}_{i}(x)=\mathcal{D}_{u}(z_{u_{i}})(x)=f_{\theta_{u},z_{u_{i}}}(x),\quad z_{u_{i}}=\{(p_{j},[c_{j};\mu])\}_{j=1}^{N_{lat}}.$ (2)
@@
Training proceeds by minimizing reconstruction losses for both input and output fields, as illustrated in Figure 1:
$\mathcal{L}^{a}=\sum_{i=1}^{n_{tr}}\frac{1}{|\mathcal{X}_{i}|}\sum_{x\in\mathcal{X}_{i}}\|a_{i}(x)-\mathcal{D}_{a}(\mathcal{E}_{a}(a_{i}))(x)\|^{2},\quad\mathcal{L}^{u}=\sum_{i=1}^{n_{tr}}\frac{1}{|\mathcal{X}_{i}|}\sum_{x\in\mathcal{X}_{i}}\|u_{i}(x)-\mathcal{D}_{u}(z_{u_{i}})(x)\|^{2}$ (3)
where $\mathcal{D}_{a}$ represents the input decoder that reconstructs input functions from their latent representations.
Input Encoder
Our encoder learns spatially anchored latent points that reconstruct the input function (geometry) via equivariant cross attention [15]:
$f_{\theta_{a},z_{a}}(x_{m})=W_{o}\sum_{j=1}^{N_{lat}}\text{att}_{m,j}\cdot v(b_{j,m},c_{j}),\quad W_{o}\in\mathbb{R}^{n_{a}\times d_{v}}$ (4)
where $b_{j,m}=\gamma(x_{m}-p_{j})$ represents the vector offset between query point $x_{m}$ and latent position $p_{j}$, composed with a Fourier feature encoding layer $\gamma$, used to overcome spectral bias and capture high frequency components [19]. The attention weights incorporate both learned affinities and spatial locality as:
$\text{att}_{m,j}=\frac{\exp\left(\frac{q(b_{j,m})^{T}k(c_{j})}{\sqrt{d_{k}}}-\sigma\|x_{m}-p_{j}\|^{2}\right)}{\sum_{\ell=1}^{N_{lat}}\exp\left(\frac{q(b_{\ell,m})^{T}k(c_{\ell})}{\sqrt{d_{k}}}-\sigma\|x_{m}-p_{\ell}\|^{2}\right)}$ (5)
where the Gaussian penalty term (controlled by window parameter $\sigma$) enhances local conditioning by penalizing distance from latent positions. The query and key functions are parameterized as: $q(b_{j,m})=W_{q}b_{j,m},k(c_{j})=W_{k}c_{j}$ $(W_{q}\in\mathbb{R}^{d_{k}\times d_{\gamma}},W_{k}\in\mathbb{R}^{d_{k}\times l_{d}})$, and the value function combines relative position information with latent features through scale and shift modulation:
$v(b_{j,m},c_{j})=(W_{v}c_{j})\odot(W_{s}b_{j,m})+W_{b}b_{j,m}$ (6)
where $W_{v}\in\mathbb{R}^{d_{v}\times l_{d}},W_{s},W_{b}\in\mathbb{R}^{d_{v}\times d_{\gamma}}$. For each input function $a_{i}$, the encoding follows a meta-learning optimization procedure based on CAVIA [23, 22]. Latent features $c$ ($\forall j\in(1,N_{lat})$) are optimized through $K=3$ gradient steps with learning rate $\alpha$:
$c_{j}^{(0)}$ $=0\quad p_{j}^{(k)}=p_{j}^{(0)}=p_{j}$ (7)
$c_{j}^{(k+1)}$ $=c_{j}^{(k)}-\alpha\nabla_{c_{j}^{(k)}}\mathcal{L}^{a}(f_{\theta_{a},c_{j}^{(k)}},a_{i})\quad\text{for }0\leq k\leq K-1$ (8)
this defines the encoder $\mathcal{E}_{a}(a_{i})=z_{a_{i}}$ as an optimization procedure. The latent positions $p_{j}$ are initialized to uniformly cover the input domain and remain fixed throughout optimization, as updating their values results in less stable training. This design choice prioritizes training stability while maintaining the spatial grounding essential for geometric representation.
During training, this inner loop optimization is performed in parallel with the global parameters $\theta_{a}$ optimization (outer loop). At inference, the global parameters are fixed, and encoding a new geometry requires only running the inner loop for a few gradient descent steps ($K$).
Output Decoder
The output decoder $\mathcal{D}_{\text{u}}$ operates on the concatenation of input latent representations $c^{a_{i}}$ and global parameters (when present) $z_{u_{i}}=\{(p_{j},[c_{j}^{a_{i}};\mu])\}_{j=1}^{N_{lat}}$ through two sequential operations: (i) latent self-attention and (ii) equivariant decoding.
The first operation enables communication between latent features, allowing mixing of global and local information. This key architectural feature operates on the latent features : $\tilde{c}_{j}=c_{j}+\operatorname*{SelfAttn}_{j}(\{c_{\ell}\}_{\ell=1}^{N_{\text{lat}}}).$ The self-attention mechanism is implemented as:
$\operatorname*{SelfAttn}_{j}(\{c_{\ell}\})\;=\;\sum_{\ell=1}^{N_{\text{lat}}}\operatorname*{softmax}_{\ell}\Bigl(\tfrac{(W_{q}c_{j})^{\top}(W_{k}c_{\ell})}{\sqrt{d_{k}}}\Bigr)(W_{v}c_{\ell}).$
This communication operation enables global feature mixing with computational complexity $O(N_{lat}^{2})$, which is typically much smaller than the spatial mesh resolution, allowing the model to capture long-range dependencies and non-local physical phenomena without the computational overhead of full spatial attention.
Following feature communication, the enhanced latent representations $\tilde{z}_{u_{i}}=\{(p_{j},\tilde{c}_{j})\}_{j=1}^{N_{lat}}$ undergo equivariant decoding using the same translation equivariant architecture as the input encoder. The output reconstruction follows the same formulation as Equation 4, with enhanced features:
$\hat{u}_{i}(x)=f_{\theta_{u},\tilde{z}_{u_{i}}}(x)=W_{o}^{\prime}\sum_{j=1}^{N_{lat}}\text{att}_{x,j}\cdot v^{\prime}(b_{x,j},\tilde{c}_{j}),\quad W_{o}^{\prime}\in\mathbb{R}^{n_{u}\times d_{v}}$
where the attention weights and value function follow the same parameterization as Equations 5-6, substituting $\tilde{c}_{j}$ for the enhanced latent features. The relative position encoding $b_{x,j}=\gamma(x-p_{j})$ maintains translation equivariance, while the spatial positions $p_{j}$ remain unchanged to preserve geometric grounding.